OpenStudy (anonymous):

Determine the equation of the line, in slope-intercept form, that is parallel to 4x-2y=-8 and contains (-8,2)

OpenStudy (anonymous):

A. y= -2x+18 B. y= 2x+18 C. y=1/2x-1/18 D. y=2x-1/18

OpenStudy (anonymous):

y=mx+b where m is the slope and b is the y-intercept Isolate the y in the equation of 4x-2y=-8 and then you are able to get the slope (m). This slope (m) is also the slope of the line parallel to the first line as parallel lines have an identical slope. 4x-2y=-8 -2y=8-4x y=(8-4x)/-2 y=-4+2x y=2x-4 The slope (m) is 2 Thus, the parallel line to the first must be y=2x+b Now, as the second line must pass through (-8,2), plug these values into y=2x+b to find the b value (y-intercept). 2=2(-8)+b 2=-16+b 18=b b=18 Finally, now that you know both the slope (m) and y-intercept (b), you cna put it together into y=mx+b y=2x+18

OpenStudy (anonymous):

so that would make it B. thanks love <3

OpenStudy (anonymous):

Yeah. You're welcome